Calculate the resistance of 110-V light bulbs rated at 25, 60, 75 and 100W.


1
Expert's answer
2021-06-28T03:32:01-0400

Gives

V=110V


"P_1=25\Omega,P_2=60\Omega,P_3=75\Omega,P_4=100\Omega"

We know that

"R=\frac{V^2}{P}"

"R_1=\frac{V^2}{P_1}=\frac{110^2}{25}=484\Omega"

"R_2=\frac{V^2}{P_2}=\frac{110^2}{60}=201\Omega"

"R_3=\frac{V^2}{P_3}=\frac{110^2}{75}=161\Omega"

"R_4=\frac{V^2}{P_4}=\frac{110^2}{100}=121\Omega"